OpenStudy (anonymous):

An aircraft is flying horizontally at a constant height of 4000 ft above a fixed observation point P. At a certain instant the angle of elevation theta is 30 degrees and decreasing, and the speed of the aircraft is 300 mi/hr. A) How fast is theta decreasing at this instant? B) How fast is the distance between the aircraft and the observation point changing at this instant?

OpenStudy (anonymous):

How do you do b?

OpenStudy (anonymous):

For part a I got -11/ 400

OpenStudy (anonymous):

You need to differentiate implicitly to find d(theta)/dt and dh/dt.

OpenStudy (anonymous):

For d(theta)/dt, I got -11/400

OpenStudy (anonymous):

I already did part A, I just need help on part B, please

OpenStudy (psymon):

Is that correct or is that just what you got?

OpenStudy (anonymous):

@Psymon -11/ 400 is what I got for part A. It could also be -0.0275 rad/ sec. I think this is correct

OpenStudy (psymon):

Nvm.

OpenStudy (anonymous):

@Psymon I found the question online and they got that answer

OpenStudy (psymon):

That was dumb, lol x_x

OpenStudy (anonymous):

What was dumb?

OpenStudy (psymon):

My suggestion for part B, lol. But yeah, Im not going to be able to get part A, not sure how ya gotit. Part B I might have. Just dont know if its right.

OpenStudy (anonymous):

I did part A like this: tan(theta)=4000/x tan(theta)=4000X^-1 do deriv (sec^2 theta)(dtheta/dt)=-4000x^-2 (dx/dt) (4/3)(d theta/dt)=-4000(sqrt{3))^-2 (440) d theta/dt=-11/400\

OpenStudy (psymon):

Im not sure where you even got sqrt(3) and the 440 from.

OpenStudy (anonymous):

You turn 300 mi/h into 440 ft/sec so that the units are the same because of the height 4000 ft

OpenStudy (psymon):

Yeah, my mind wouldve never caught that. Alright, I wont bother you with my part B then.

OpenStudy (anonymous):

Noobs xD

OpenStudy (anonymous):

A solution using Mathematica for the calculations is attached.